# X-Reflect: Cross-Reflection Prompting for Multimodal Recommendation

### v1

@@ Abstract
Large Language Models (LLMs) and Large Multimodal Models (LMMs) have been shown to enhance the effectiveness of enriching item descriptions, thereby improving the accuracy of recommendation systems. However, most existing approaches either rely on text-only prompting or employ basic multimodal strategies that do not fully exploit the complementary information available from both textual and visual modalities. This paper introduces a novel framework, Cross-Reflection Prompting, termed X-Reflect, designed to address these limitations by prompting LMMs to explicitly identify and reconcile supportive and conflicting information between text and images. By capturing nuanced insights from both modalities, this approach generates more comprehensive and contextually richer item representations. Extensive experiments conducted on two widely used benchmarks demonstrate that our method outperforms existing prompting baselines in downstream recommendation accuracy. Additionally, we evaluate the generalizability of our framework across different LMM backbones and the robustness of the prompting strategies, offering insights for optimization. This work underscores the importance of integrating multimodal information and presents a novel solution for improving item understanding in multimodal recommendation systems.
Introduction
@@
Leveraging Large Language Models (LLMs) or Large Multimodal Models (LMMs) to enrich item descriptions has been shown to enhance the effectiveness of recommendation systems (Lyu et al. 2024; Xi et al. 2023; Li et al. 2023). By generating detailed and contextually relevant descriptions, LLMs and LMMs contribute to more accurate and personalized recommendations. However, many existing methods either rely on text-only prompting strategies or use basic multimodal prompting techniques that do not fully utilize the information available from both textual and visual modalities (Liu et al. 2024c; Baltrušaitis, Ahuja, and Morency 2018).
@@
For instance, a movie described textually as a “light-hearted romantic comedy” may have a poster that, with its dark tones and serious expressions, suggests underlying tension. A text-only approach might overlook these visual cues, while a simple multimodal strategy might superficially combine the text and image without resolving their differences, leading to an incomplete or misleading representation of the item. Similarly, a grocery item labeled as “organic and healthy” might have packaging that subtly suggests indulgence, like rich colors or gourmet styling. While neither the text nor the image alone may convey the full character of the product, analyzing them together can highlight these additional dimensions, such as the balance between health and indulgence. These examples illustrate the limitations of current approaches and the need for more sophisticated methods that can effectively integrate and interpret information across modalities.
@@
To recommend items more effectively, a deeper understanding of these items is necessary, which cannot be achieved by neglecting the rich information provided by additional modalities, such as images. Visual content often captures aspects that text alone cannot fully convey, including aesthetic qualities, emotional tone, or specific details that may significantly influence user preferences (Huang et al. 2019). Therefore, it is critical to develop methods that can seamlessly integrate and reconcile information from both text and images.
@@
To address this need, we introduce a novel Cross-Reflection Prompting framework termed X-Reflect. This method, as illustrated in Figure 1, prompts LMMs to process textual and visual information concurrently, explicitly identifying and reconciling any supportive or conflicting elements between these modalities. By capturing the nuanced information present in both text and images, cross-reflection prompting enables the generation of more diverse and contextually rich item representations.
@@
We conduct extensive experiments on two widely used recommendation benchmarks, demonstrating that our framework consistently outperforms baseline prompting methods in recommendation accuracy.
We assess the generalizability of our framework by applying it to multiple LMM backbones, showing that it remains effective across different model architectures.
In summary, this work underscores the importance of fully exploiting multimodal information in recommendation systems and presents a novel framework that enhances item understanding through a more integrated analysis of textual and visual data.
Related Work
Multimodal Recommendation System
Traditional unimodal recommendation systems, which depend on a single data source such as textual descriptions, often encounter challenges like data sparsity, cold start problems, noise, and redundancy. To mitigate these issues, researchers have developed multimodal recommendation systems that integrate diverse types of data, including text, images, audio, and video. These systems enable a more comprehensive understanding of content, as different modalities can offer varied perspectives and insights. Moreover, when one modality suffers from poor quality or sparse data, other modalities can compensate, enhancing the system’s overall robustness and reliability (Wei et al. 2019; Lyu and Luo 2022; Wang, Chen, and Zhu 2021; Huang et al. 2019; Tao et al. 2020).
@@
Prompting Strategies for Recommendation
@@
There has been limited exploration of multimodal prompting strategies in the context of recommendation. One notable exception is Rec-GPT4V (Liu et al. 2024c), which introduced a reasoning scheme called Visual-Summary Thought (VST). This method leverages large vision-language models (LVLMs) for multimodal recommendation by using user history as in-context user preferences and generating item image summaries to enhance the recommendation process.
@@
Cross-Reflection Prompting
@@
This paper aims to design a framework that can more effectively leverage Large Multimodal Models to enrich item descriptions by fully exploiting information from both textual and visual modalities. To achieve this, we propose a novel approach called Cross-Reflection Prompting, referred to as X-Reflect.
The core idea of cross-reflection prompting is to provide LMMs with both textual and visual information as context, while explicitly instructing the models to identify and reconcile whether the textual and visual information supports or contradicts each other. The responses generated by the LMMs are then converted into embeddings, which serve as input item embeddings for the subsequent recommendation module, as illustrated in Figure 2.
We present two distinct strategies for implementing cross-reflection prompting: separate prompting and combined prompting.
Separate Prompting
In separate prompting, we first prompt two models, and to independently augment the text input T and image input I, respectively:
where and represent the intermediate generated responses. Next, another model, , is employed to generate a more comprehensive response based on these intermediate responses. In the prompt for , cross-reflection is incorporated as follows:
@@
Combined Prompting
In combined prompting, we use a single model, , to generate the response based on the text and image associated with the item through cross-reflection prompting:
“Please provide a summary of the item characteristics based on the provided text description and image. The item description is as follows: {text description}. Evaluate whether the information presented in the text and the image are supportive of each other or if there are any discrepancies. If any conflicts are identified, please address these conflicts.”
In both prompts, the parts related to cross-reflection are highlighted in purple. Inspired by the findings of Lyu et al. (2024), we concatenate the embeddings of these responses to construct item representations, which will be used in the subsequent recommendation module.
Remark. Similar to prior augmentation methods, cross-reflection prompting can be considered a plug-and-play module in the recommendation system, serving solely to augment item-side input, thus providing significant flexibility for the recommendation module. Moreover, although Figure 2 indicates all models as LMMs, only and need to be LMMs. and can be Large Language Models that do not take images as input. can be omitted, allowing the use of the original item description directly. This further enhances the flexibility of the entire framework. Lastly, both proprietary and open-source models can be used. We leave the investigation of incorporating supervised fine-tuning (SFT) and continuous pre-training of the models to improve the quality of responses for future work.
Experiment Setup
Datasets
Dataset | # Interactions | # Items | # Users |
MovieLens-1M | 1,000,209 | 3,706 | 6,040 |
Amazon-Software | 6,251 | 1,729 | 1,004 |
We use two widely recognized recommendation benchmarks: MovieLens-1M (Harper and Konstan 2015) and Amazon-Software (Ni, Li, and McAuley 2019). The dataset statistics are provided in Table 1.
MovieLens-1M: This dataset contains anonymous user ratings collected from MovieLens. The movie poster URLs are obtained from the movielens-poster repository.111https://github.com/babu-thomas/movielens-posters
@@
LLM-Rec (Lyu et al. 2024): This work proposes four distinct prompting strategies to enrich item descriptions. We use the recommendation-driven strategy as our baseline. The exact prompt is: “The description of an item is as follows: {item description}, what else should I say if I want to recommend it to others?”
@@
Rec-GPT4V (Liu et al. 2024c): This baseline prompts LMMs to generate item image summaries and integrates image comprehension in the natural language space with item titles to query user preferences. For fair comparison, we only use the instruction to generate item image summaries. The exact prompt is “What’s in this image?”
CoT (Wei et al. 2022): The Chain-of-Thought (CoT) prompting guides LLMs to generate step-by-step reasoning before delivering a final answer. Many researchers have explored its adaptation into the multimodal context (Zhang et al. 2024; Zheng et al. 2023). In our implementation, the prompt is “The item description is as follows: {item description}. Please think step by step and describe this item based on both the description and the input image. Your response should highlight how combining both the text and image provides a comprehensive understanding of the item.”
Visual-only prompting: This method prompts an LMM to describe the image: “Describe the image.”
Visual-textual prompting: It prompts an LMM to describe the item based on both the image and the original item description: “The item description is as follows: {item description}. Please describe this item based on both the description and the input image. Your response should highlight how combining both the text and image provides a comprehensive understanding of the item.”
@@
Text + Image: This baseline uses the embeddings of the original item description combined with CLIP embeddings (Radford et al. 2021) of the associated image as input to the recommendation module.
Evaluation Protocols
@@
Large Multimodal Models
For fair comparison, we use the same language model across all prompting strategies. For X-Reflect and other multimodal prompting baselines, we use GPT-4V (OpenAI 2023) (gpt-4-vision-preview variant). To test the effectiveness and generalizability of our framework, we also evaluate it using two open-source LMMs: LLaVA-1.5 (Liu et al. 2024a) (llava-1.5-7b-hf), an enhanced version of LLaVa (Liu et al. 2024b), and InternLM-XComposer2 (Dong et al. 2024) (internlm-xcomposer2-vl-7b), a vision-language large model based on InternLM2 (Cai et al. 2024). For text-only prompting baselines, we use GPT-4 (gpt-4 variant) (OpenAI 2023).
Implementation Details
Recommendation Module
Following Lyu et al. (2024), we evaluate different prompting baselines using a two-tower model (Wang, Zhu, and He 2021), designed to predict the likelihood of a user interacting with an item. The model consists of two towers that process user and item features into dense embeddings. The interaction likelihood is computed using the dot product of these embeddings. We use Sentence-BERT (Reimers and Gurevych 2019) to generate textual embeddings from responses, specifically using the all-MiniLM-L6-v2 model variant. These item embeddings serve as input to the item tower. The text encoder remains frozen to minimize computational costs. For the user tower, we employ an embedding table to convert user IDs into latent representations, with the output dimension set to 128 for both datasets.
Model Training
@@
Grid Search for Hyperparameters
Model parameters are initialized randomly following a Gaussian distribution. We optimize the framework using the AdamW algorithm (Loshchilov and Hutter 2017) with a weight decay of 0.0005. Hyperparameter grids for learning rate and dropout are set to and , respectively. Early stopping is implemented with a window size and evaluation frequency of 5. The batch size is set to 4096 across all baselines. Hyperparameters that yield the highest Recall@K on the validation set are used for final testing.
Due to computational constraints, we do not perform grid search for LMM or LLM hyperparameters, instead following the settings from Lyu et al. (2024).
Computing Resources
@@
Results
Comparison with Baseline Prompting Strategies
@@ Results
First, prompting methods that integrate both text and image generally outperform those relying solely on text. However, on the Amazon-Software dataset, the text-only prompting strategy LLM-Rec (Lyu et al. 2024) shows notably strong performance. The specific LLM-Rec variant implemented in this work involves instructing the LLM to describe the item with a focus on recommendation purposes. Upon closer examination, it appears that the original Amazon-Software descriptions, which contain highly technical details, may not generalize well. The responses generated through LLM-Rec, which tend to be less technical, likely contribute to its superior performance.
Second, among the prompting methods that combine text and image, those incorporating cross-reflection consistently outperform those that do not take cross-reflection into account.
Third, similar to the findings in LLM-Rec, our method, which leverages augmented information from the cross-reflection prompting components, achieves superior performance compared to all other baselines.
MovieLens-1M Amazon-Software Precision@10 Recall@10 NDCG@10 Precision@10 Recall@10 NDCG@10 Item Popularity 0.0426 0.0428 0.0530 0.0082 0.0777 0.0571 Text 0.2877 0.2424 0.3597 0.0231 0.2175 0.1568 KAR (Xi et al. 2023) 0.2990 0.2575 0.3724 0.0189 0.1686 0.1175 LLM-Rec (Lyu et al. 2024) 0.2965 0.2547 0.3710 0.0424 0.3911 0.2843 Text + Image 0.2936 0.2529 0.3645 0.0268 0.2522 0.1966 Rec-GPT4V (Liu et al. 2024c) 0.3004 0.2589 0.3748 0.0254 0.2401 0.2000 CoT (Wei et al. 2022) 0.3039 0.2636 0.3806 0.0305 0.2859 0.2157 Visual-only prompting 0.3025 0.2626 0.3770 0.0274 0.2591 0.1988 Visual-textual prompting 0.3036 0.2633 0.3790 0.0267 0.2490 0.1914 Cross-reflection - separate 0.3060 0.2649 0.3830 0.0385 0.3537 0.2661 Cross-reflection - combined 0.3067 0.2681 0.3822 0.0318 0.2968 0.2414 X-Reflect 0.3111 0.2746 0.3900 0.0431 0.3969 0.3024 (+5.96) (+8.58) (+7.00) (+60.82) (+57.38) (+53.81)
Top-1 Recommendation Performance
MovieLens-1M Amazon-Software Precision@1 Recall@1 Precision@1 Recall@1 Text 0.4109 0.0401 0.1050 0.1017 KAR (Xi et al. 2023) 0.4248 0.0431 0.2108 0.2039 LLM-Rec (Lyu et al. 2024) 0.4150 0.0412 0.2191 0.2135 Text + Image 0.3966 0.0390 0.1550 0.1508 Rec-GPT4V (Liu et al. 2024c) 0.4115 0.0415 0.1863 0.1826 CoT (Wei et al. 2022) 0.4214 0.0430 0.1803 0.1744 Visual prompting 0.4103 0.0407 0.1892 0.1858 Visual-textual prompting 0.4131 0.0415 0.1813 0.1766 Cross-reflection - separate 0.4226 0.0424 0.2118 0.2034 Cross-reflection - combined 0.4119 0.0413 0.2080 0.2027 X-Reflect 0.4232 0.0436 0.2271 0.2177 (+6.71) (+11.79) (+46.52) (+44.36)
Table 3 displays the recommendation performance for Top-1 recommendations, measured by Precision@1 and Recall@1. These metrics assess the quality of recommendations when only one item is suggested. Since only one item is recommended, NDCG@1 simplifies to Precision@1. X-Reflect demonstrates strong performance in Top-1 recommendations, achieving the highest Precision@1 and Recall@1 on the Amazon-Software dataset, and the highest Recall@1 with the second-highest Precision@1 on MovieLens-1M. KAR (Xi et al. 2023) achieves the highest Precision@1 on MovieLens-1M, likely because it uses open-world knowledge to enrich item descriptions. For movies, unique characteristics that are better represented as knowledge may allow KAR to precisely augment special information, making it particularly effective in Top-1 recommendations. However, as seen in the Top-10 recommendation results, KAR’s performance diminishes, potentially due to its limited ability to enrich general information effectively.
Generalizability of the Proposed Framework
To assess the generalizability of our framework, we replace the GPT-4V backbone with two other widely used LMMs–InternLM and LLaVA–while keeping the experimental settings consistent across all baselines. Tables 5 and 4 show the corresponding recommendation performance. With LLaVA as the backbone, X-Reflect continues to outperform other prompting baselines. When InternLM is used as the backbone, our method remains one of the top-performing frameworks, achieving the highest Precision@10 across both datasets and excelling on the Amazon-Software dataset.
MovieLens-1M Amazon-Software Precision@10 Recall@10 NDCG@10 Precision@10 Recall@10 NDCG@10 Text + Image 0.2936 0.2529 0.3645 0.0268 0.2522 0.1966 Rec-GPT4V (Liu et al. 2024c) 0.2946 0.2523 0.3666 0.0309 0.2916 0.2245 CoT (Wei et al. 2022) 0.3001 0.2603 0.3746 0.0255 0.2406 0.2033 Visual-only prompting 0.2958 0.2542 0.3689 0.0313 0.2939 0.2388 Visual-textual prompting 0.2998 0.2570 0.3740 0.0317 0.2962 0.2380 Cross-reflection - separate 0.2909 0.2464 0.3635 0.0438 0.4048 0.3257 Cross-reflection - combined 0.3002 0.2573 0.3752 0.0278 0.2602 0.2210 X-Reflect 0.3002 0.2591 0.3736 0.0457 0.4172 0.3232 (+2.25) (+2.45) (+2.50) (+70.52) (+65.42) (+64.39)
MovieLens-1M Amazon-Software Precision@10 Recall@10 NDCG@10 Precision@10 Recall@10 NDCG@10 Text + Image 0.2936 0.2529 0.3645 0.0268 0.2522 0.1966 Rec-GPT4V (Liu et al. 2024c) 0.3002 0.2569 0.3756 0.0320 0.3039 0.2481 CoT (Wei et al. 2022) 0.3007 0.2574 0.3760 0.0288 0.2702 0.2106 Visual-only prompting 0.2968 0.2544 0.3700 0.0292 0.2747 0.2184 Visual-textual prompting 0.3005 0.2578 0.3754 0.0309 0.2884 0.2341 Cross-reflection - separate 0.3017 0.2609 0.3765 0.0398 0.3696 0.2970 Cross-reflection - combined 0.2976 0.2558 0.3715 0.0317 0.2969 0.2473 X-Reflect 0.3026 0.2640 0.3783 0.0411 0.3829 0.3049 (+3.07) (+4.39) (+3.79) (+53.36) (+51.82) (+55.09)
Further Discussion on Augmented Content
To better understand the impact of cross-reflection, we compute the cosine similarity between the embeddings of the original text description and the image description, as well as the responses generated through text-only prompting (LLM-Rec), multimodal prompting (CoT), and cross-reflection (separate) multimodal prompting.
For each item in the MovieLens-1M dataset, we convert its text description and image description into embeddings, defining the cosine similarity between these two embeddings as reference similarity. A higher value indicates greater semantic similarity between the text and image descriptions, serving as a proxy for how much additional information the image description contains compared to the original text description.
Similarly, we compute the cosine similarity between the text description embeddings and the responses generated through text-only prompting, multimodal prompting, and cross-reflection multimodal prompting. We then compare these cosine similarities with the reference similarity. The results are presented in Figure 3.
In each subplot, the x-axis represents the cosine similarity between the embeddings of the text and image descriptions, while the y-axis shows the cosine similarity between the embeddings of the text description and the responses generated through the other three prompting baselines.
The key finding is that cross-reflection multimodal prompting generates semantically more diverse information compared to other prompting strategies and image descriptions. This aligns with our earlier discussion, where we note that visual information from item images contains additional details not present in the original text description. However, directly prompting LMMs to describe an image may not elicit a comprehensive understanding of the item.
Even combining text and image in the prompt is not sufficient; explicit instruction is necessary to direct LMMs to focus on non-overlapping information across modalities. Although not identical, a similar pattern has been observed in multi-input image prompting (Zhang et al. 2024), where explicit instructions to focus on similarities and differences among multiple image inputs improve multimodal reasoning in LMMs.
It is important to note that the utility of augmented information in downstream recommendation tasks is still under investigation. While this experiment demonstrates that semanticallt more diverse information often leads to better performance, this is not an absolute rule. As seen in the experiments, generating less granular responses can sometimes also enhance performance in downstream tasks. This remains an ongoing area of research in the recommendation domain, and we plan to explore this further in future work.
Ablation on Prompts
To assess the robustness of our prompting framework, we modify our prompts to instruct LMMs to generate only unique keywords while maintaining the explicit instruction to identify whether the text and image information is supportive or contradictory. The results are presented in Table 6. This variant, termed X-Reflect-keyword, shows that X-Reflect is robust across different prompt designs. The performance remains consistent with the inclusion of explicit cross-reflection instructions.

### v2

@@ Abstract
Large Language Models (LLMs) have been shown to enhance the effectiveness of enriching item descriptions, thereby improving the accuracy of recommendation systems. However, most existing approaches either rely on text-only prompting or employ basic multimodal strategies that do not fully exploit the complementary information available from both textual and visual modalities. This paper introduces a novel framework, Cross-Reflection Prompting, termed X-Reflect, designed to address these limitations by prompting Multimodal Large Language Models (MLLMs) to explicitly identify and reconcile supportive and conflicting information between text and images. By capturing nuanced insights from both modalities, this approach generates more comprehensive and contextually rich item representations. Extensive experiments conducted on two widely used benchmarks demonstrate that our method outperforms existing prompting baselines in downstream recommendation accuracy. Furthermore, we identify a U-shaped relationship between text–image dissimilarity and recommendation performance, suggesting the benefit of applying multimodal prompting selectively. To support efficient real-time inference, we also introduce X-Reflect-keyword, a lightweight variant that summarizes image content using keywords and replaces the base model with a smaller backbone, achieving nearly 50% reduction in input length while maintaining competitive performance. This work underscores the importance of integrating multimodal information and presents an effective solution for improving item understanding in multimodal recommendation systems.
[…]
X-Reflect: Cross-Reflection Prompting for Multimodal Recommendation
Hanjia Lyu1, Ryan Rossi2, Xiang Chen2, Md Mehrab Tanjim2, Stefano Petrangeli2, Somdeb Sarkhel2, Jiebo Luo1 1University of Rochester, 2Adobe Research hlyu5@ur.rochester.edu, jluo@cs.rochester.edu
1 Introduction
@@
Original item descriptions often lack sufficient detail or context, which limits the ability of recommendation models to effectively learn user preferences. Leveraging Large Language Models (LLMs) to enrich item descriptions has been shown to enhance the effectiveness of recommendation systems Lyu et al. (2024); Xi et al. (2023); Li et al. (2023). By generating richer, more informative, and contextually relevant content, LLMs enable the model to better capture item semantics and align them with user interests, leading to more accurate and personalized recommendations. However, many existing methods either rely on text-only prompting strategies or use basic multimodal prompting techniques that do not fully utilize the information available from both textual and visual modalities Liu et al. (2024c); Baltrušaitis et al. (2018).
@@
For instance, a movie described textually as a “light-hearted romantic comedy” may have a poster that, with its dark tones and serious expressions, suggests underlying tension. A text-only approach might overlook these visual cues, while a simple multimodal strategy might superficially combine the text and image without resolving their differences, leading to an incomplete or misleading representation of the item. Similarly, a grocery item labeled as “organic and healthy” might have packaging that subtly suggests indulgence, like rich colors or gourmet styling. While neither the text nor the image alone may convey the full character of the product, analyzing them together can highlight these additional dimensions, such as the balance between health and indulgence. These examples illustrate the limitations of current approaches and the need for more sophisticated methods that can effectively integrate information across modalities.
@@
To recommend items more effectively, a deeper understanding of these items is necessary, which cannot be achieved by neglecting the rich information provided by additional modalities, such as images. Visual content often captures aspects that text alone cannot fully convey, including aesthetic qualities, emotional tone, or specific details that may significantly influence user preferences Huang et al. (2019). However, merely describing the visual content may not be optimal. As shown in Fig. 1, cross-modal information is most beneficial in multimodal prompting when there is a moderate level of misalignment between text and images. At both extremes—where text and images are either highly similar (left) or highly dissimilar (right)—while both approaches improve recommendation performance compared to using only textual information, their effectiveness is lower than in cases of moderate misalignment. Therefore, it is critical to develop methods that can seamlessly integrate and reconcile information from both text and images.
@@
To address this need, we introduce a novel Cross-Reflection Prompting framework termed X-Reflect. This method, as illustrated in Fig. 2, prompts Multimodal Large Language Models (MLLMs) to process textual and visual information concurrently, explicitly identifying and reconciling any supportive or conflicting elements between these modalities. By capturing the nuanced information present in both text and images, cross-reflection prompting enables the generation of more diverse and contextually rich item representations.
@@
We conduct extensive experiments on two widely used recommendation benchmarks, demonstrating that our framework consistently outperforms baseline prompting methods across different MLLM backbones.
Through qualitative and quantitative analysis, we show that X-Reflect provides more insightful and discrepancy-aware outputs compared to other multimodal prompting methods such as Rec-GPT4V and Chain-of-Thought prompting.
We uncover a U-shaped relationship between text–image dissimilarity and recommendation performance, suggesting that multimodal prompting is most beneficial when the modalities provide moderately dissimilar, complementary information. To support real-time recommendation scenarios, we introduce X-Reflect-keyword, a latency-efficient variant that summarizes image content using keywords and utilizes a smaller MLLM backbone, achieving nearly 50% token reduction while maintaining competitive performance.
This work underscores the importance of fully exploiting multimodal information in recommendation systems and presents a novel framework that enhances item understanding through a more integrated analysis of textual and visual data.
2 Related Work
2.1 Multimodal Recommendation System
Traditional unimodal recommendation systems, which depend on a single data source such as textual descriptions, often encounter challenges like data sparsity, cold start problems, noise, and redundancy. To mitigate these issues, researchers have developed multimodal recommendation systems that integrate diverse types of data, including text, images, audio, and video. These systems enable a more comprehensive understanding of content, as different modalities can offer varied perspectives and insights. Moreover, when one modality suffers from poor quality or sparse data, other modalities can compensate, enhancing the system’s overall robustness and reliability Wei et al. (2019); Lyu and Luo (2022); Wang et al. (2021b); Huang et al. (2019); Tao et al. (2020).
@@ 2.1 Multimodal Recommendation System
2.2 Prompting Strategies for Recommendation
@@
There has been limited exploration of multimodal prompting strategies in the context of recommendation. One notable exception is Rec-GPT4V Liu et al. (2024c), which introduced a reasoning scheme called Visual-Summary Thought (VST). This method leverages large vision-language models (LVLMs) for multimodal recommendation by using user history as in-context user preferences and generating item image summaries to enhance the recommendation process.
@@
3 Cross-Reflection Prompting
@@
This paper aims to design a framework that can more effectively leverage Multimodal Large Language Models to enrich item descriptions by fully exploiting information from both textual and visual modalities. To achieve this, we propose a novel approach called Cross-reflection Prompting, referred to as X-Reflect.
The core idea of cross-reflection prompting is to provide MLLMs with both textual and visual information as context, while explicitly instructing the models to identify and reconcile whether the textual and visual information supports or contradicts each other. The responses generated by the MLLMs are then converted into embeddings, which serve as input item embeddings for the subsequent recommendation module.
We first prompt the model twice to augment the text input T and image input I:
where and represent the intermediate generated responses. Following Lyu et al. (2024), the prompt for the text-only input is:
“The description of an item is as follows: {T}. What else should I say if I want to recommend it to others?”
The prompt for image-only input is:
“Describe the image.”
Next, the same model, , is employed again to generate a more comprehensive response based on these intermediate responses. In the final prompt, cross-reflection is incorporated as follows:
@@
Remark. Similar to prior augmentation methods, cross-reflection prompting can be considered a plug-and-play module in the recommendation system, serving solely to augment item-side input, thus providing significant flexibility for the recommendation module. We leave the investigation of incorporating supervised fine-tuning (SFT) and continuous pre-training of the models to improve the quality of responses for future work.
4 Experiment Setup
To evaluate the effectiveness of the proposed prompting strategy, we assess how well the recommendation model performs when using the augmented item descriptions. Specifically, we convert both the original item descriptions and the augmented responses into embeddings, then concatenate them to form the input to the recommendation model. Improvements in recommendation performance indicate the value added by the prompting strategy.
4.1 Datasets
Dataset # Interactions # Items # Users MovieLens-1M 1,000,209 3,706 6,040 Amazon-Software 6,251 1,729 1,004
We use two widely recognized recommendation benchmarks: MovieLens-1M Harper and Konstan (2015) and Amazon-Software Ni et al. (2019). The dataset statistics are provided in Table 1.
MovieLens-1M: This dataset contains anonymous user ratings collected from MovieLens. The movie poster URLs are obtained from the movielens-poster repository.111License: https://files.grouplens.org/datasets/movielens/ml-1m-README.txt
@@ 4.2 Baselines
LLM-Rec Lyu et al. (2024): This work proposes four distinct prompting strategies to enrich item descriptions. We use the recommendation-driven strategy as our baseline. The exact prompt is:
“The description of an item is as follows: {item description}, what else should I say if I want to recommend it to others?”
@@
Rec-GPT4V Liu et al. (2024c): This baseline prompts MLLMs to generate item image summaries and integrates image comprehension in the natural language space with item titles to query user preferences. For fair comparison, we only use the instruction to generate item image summaries. The exact prompt is:
“What’s in this image?”
Since this instruction only produces an image description, the approach can be regarded as a visual-only prompting method.
CoT Wei et al. (2022): The Chain-of-Thought (CoT) prompting guides LLMs to generate step-by-step reasoning before delivering a final answer. Many researchers have explored its adaptation into the multimodal context Zhang et al. (2024); Zheng et al. (2023). In our implementation, the prompt is:
“The item description is as follows: {item description}. Please think step by step and describe this item based on both the description and the input image. Your response should highlight how combining both the text and image provides a comprehensive understanding of the item.”
@@
Text + Image: This baseline uses the embeddings of the original item description combined with CLIP embeddings Radford et al. (2021) of the associated image as input to the recommendation module.
4.3 Evaluation Protocols
@@
4.4 Multimodal Large Language Models
For fair comparison, we use the same language model across all prompting strategies. For X-Reflect and other multimodal prompting baselines, we use GPT-4o Hurst et al. (2024) (gpt-4o-2024-08-06). To test the effectiveness and generalizability of our framework, we also evaluate it using another open-source MLLM: LLaVA-1.5 Liu et al. (2024a) (llava-1.5-7b-hf), an enhanced version of LLaVa Liu et al. (2024b). For text-only prompting baselines, we use GPT-4o-mini (gpt-4o-mini) Hurst et al. (2024). We use the default values of the hyperparameters (e.g., temperature) for all language models.
4.5 Implementation Details
4.5.1 Recommendation Module
Following Lyu et al. (2024), we evaluate different prompting baselines using a two-tower model Wang et al. (2021a), designed to predict the likelihood of a user interacting with an item. The model consists of two towers that process user and item features into dense embeddings. The interaction likelihood is computed using the dot product of these embeddings. We use Sentence-BERT Reimers and Gurevych (2019) to generate textual embeddings from responses, specifically using the all-MiniLM-L6-v2 model variant. These item embeddings serve as input to the item tower. The text encoder remains frozen to minimize the computational costs during the training phase. For the user tower, we employ an embedding table to convert user IDs into latent representations, with the output dimension set to 128 for both datasets.
4.5.2 Model Training
@@
4.5.3 Grid Search for Hyperparameters
Model parameters are initialized randomly following a Gaussian distribution. We optimize the framework using the AdamW algorithm Loshchilov and Hutter (2017) with a weight decay of 0.0005. Hyperparameter grids for learning rate and dropout are set to and , respectively. Early stopping is implemented with a window size and evaluation frequency of 5. The batch size is set to 4096 across all baselines. Hyperparameters that yield the highest Recall@K on the validation set are used for final testing.
4.5.4 Computing Resources
@@
MovieLens-1M Amazon-Software Precision@10 Recall@10 NDCG@10 Precision@10 Recall@10 NDCG@10 Item Popularity 0.0426 0.0428 0.0530 0.0082 0.0777 0.0571 Text 0.2721 0.2452 0.3464 0.0503 0.4503 0.3278 KAR Xi et al. (2023) 0.2828 0.2600 0.3592 0.0261 0.2255 0.1553 LLM-Rec Lyu et al. (2024) 0.2827 0.2573 0.3604 0.0503 0.4483 0.3371 Text + Image 0.2781 0.2551 0.3523 0.0467 0.4182 0.2915 Rec-GPT4V Liu et al. (2024c) 0.2844 0.2618 0.3614 0.0478 0.4268 0.3174 CoT Wei et al. (2022) 0.2825 0.2596 0.3589 0.0494 0.4423 0.3218 X-Reflect 0.2847 0.2619 0.3621 0.0521 0.4654 0.3266 (+2.37) (+2.67) (+2.78) (+11.56) (+11.29) (+12.04)
5 Results
5.1 Comparison with Baseline Prompting Strategies
@@
First, prompting methods that integrate both text and image generally outperform those relying solely on text. However, on the Amazon-Software dataset, the text-only prompting strategy LLM-Rec Lyu et al. (2024) shows notably strong performance. The specific LLM-Rec variant implemented in this work involves instructing the LLM to describe the item with a focus on recommendation purposes. Upon closer examination, it appears that the original Amazon-Software descriptions, which often contain condensed phrases and are uninformative, may not generalize well. For instance, an original description such as “Nolo SBQ11R Quicken Legal Business Pro 2011” offers minimal context. In contrast, LLM-Rec generates enriched, purpose-driven text like:
“I highly recommend the Nolo SBQ11R Quicken Legal Business Pro 2011 for anyone looking to streamline their legal business operations. This software is designed specifically for legal professionals and offers a comprehensive suite of tools to manage your practice efficiently .”
Such detailed outputs likely account for LLM-Rec’s strong performance on this dataset.
Second, X-Reflect consistently outperforms other multimodal prompting methods across both datasets. When comparing the enriched texts generated by Rec-GPT4V, CoT, and X-Reflect, we find that X-Reflect provides more insightful information due to its specific instruction to reason about consistencies and discrepancies between the text and image. For instance, for one item, X-Reflect’s response notes:
“The item description states there are ‘4,500 exercises,’ whereas the image description mentions ‘800 new exercises.’ This could imply that the 800 exercises are new additions to the existing total.”
This level of reasoning is absent from the other methods: Rec-GPT4V omits any mention of exercises, and CoT merely restates the original description without identifying the discrepancy. This suggests that X-Reflect ’s cross-modal reasoning-based instruction leads to more informative and contextually grounded outputs.